If one number is 12 more than another. The sum of the smaller number and twice the larger number is 39 then what is the larger number
Answer by checkley71(8403)

X+12=Y
X+2(X+12)=39
X+2X+24=39
3X=39-24
3X=15
X=15/3
X=5 THUS
5+12=Y
17=Y
PROOF
5+2(5+12)=39
5+2*17=39
5+34=39
39=39